Please help. Any of the answer choices below each table is an answer to one of the Q’s.
oess deffective ull r 014 012 015 b. 121,675 013 016 124,000.00 c. 2,325 d. 1,538.46 e.78,461.54 a. 7.64 017 018 019 140 12 e. 0.24 a. 145,775 b. 144,294.98 c. 1,480.02 d. 5,785 effective r Ci 112.78 1,500.00 7.25% Q2 022 b. 360 c. 1,612.78 d. 70,673.18 e. 22,673.18 a. 6.66 EY b. 62,482.56 c. 24 d. 22,673.18 e.0.5 a. 64 Which one is EY APR comp to M.(in yrs) frequency 024 Cit 100,000.00 10.25% a. 139,995.19 b, 49,158.52 . 149,158.52 d. 39,995.19 e. 16
Answer #1
D
Principal APR T-t $ int FV d B-discount Proceeds
     140,000 5.50% 0.75                                   5,775           145,736 4.25%                        60         145,775
=270/360 =140000*5.5%*(270/360) 145736-140000
T to M (days T to M (yrs) t1 t2 t2-t1
270 0.75 15-Mar 15-Sep 184
=270/360
Principal APR t to M (yrs) Freq t*f $ int FV effective r
1500 7.25% 1 360 360 $112.78 $1,612.78 7.52%
48000 6.50% 6 4 24 $22,673.18 $70,673.18 6.66%
E
FV APR t to M (yrs) Freq t*f PV
     240,000 8.50% 16 4 64    62,482.56
Principal APR t to M (yrs) Freq t*f $ int FV
Citi 100,000 10.25% 4 2 8 $49,158.52 $149,158.52
Chase 100,000 8.50% 4 4 16 $39,995.19 $139,995.19
Citi is better